OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 17 and a15 = -39 ?

OpenStudy (anonymous):

@eliassaab : What is the 7th term of the geometric sequence where a1 = -4,096 and a4 = 64?

OpenStudy (anonymous):

r = 64

OpenStudy (anonymous):

\[-39=14 r+17 r=-4 \] for the first problem

OpenStudy (anonymous):

\[-39=14 r+17\\ r=-4 \]

OpenStudy (anonymous):

my choices for that one are: A) -147 B) -143 C) -139 D) -135

OpenStudy (anonymous):

a41=17+ 40 (-4)=-143

OpenStudy (anonymous):

thats what i got.

OpenStudy (anonymous):

Excellent

OpenStudy (anonymous):

A) -1 B) 4 C) 1 D) -4 options for the second one... i got D

OpenStudy (anonymous):

It is A.

OpenStudy (anonymous):

Are you sure a1 isn't 4096?

OpenStudy (anonymous):

okay, i'll work it again.

OpenStudy (anonymous):

@telliott99 i'm sure . its -

OpenStudy (anonymous):

-4096 r^3 =64 this gives r = -1/4 -4096(-1/4)^6= -1

OpenStudy (anonymous):

i got -1 :)

OpenStudy (anonymous):

Excellent

OpenStudy (anonymous):

I see, sorry to cause confusion.

OpenStudy (anonymous):

-4096 1024 -256 64 -16 4 -1